The three angles in trangle POR are in the ratio of 6:26:13. What is the measure of each angle?
Answer by checkley71(8403)

IF YOU ADD 16+26+13=55
NOW DIVIDE 180 BY 55
180/55=3.27272727 THIS IS THE FACTOR TO MULTIPLY EACH RATIO BY TO FIND THE SPECIFIC ANGLES.
16*3.272727=52.363636 DEGREES FOR ANGLE #1
26*3.272727=85.090909 DEGREES FOR ANGLE #2
13*3.272727=42.545454 DEGREES FOR ANGLE #3
NOW ADD THEM.
52.363636
85.090909
42.545454
-----------------
179.999999 OR ROUNDED TO 180 DEGREES.